OpenStudy (anonymous):

1. Let f(x) = 3x – 5. Find its inverse represented by the function notation f^ –1(x).

OpenStudy (anonymous):

what would you do if you had to solve \[3x-5=17\]?

OpenStudy (anonymous):

first you would add 5 to both sides, which would give you 3x= 22 and then you would divide 3 on both sides?

OpenStudy (anonymous):

yes and that is what the inverse does add 5, divide by 3 \[f^{-1}(x)=\frac{x+5}{3}\]

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

yes that is it

OpenStudy (anonymous):

oh cool, thanks! Could you help me with another question?

OpenStudy (anonymous):

you could write \[x=3y-5\] and sovle for \(y\) to get \[x+5=3y\] \[y=\frac{x+5}{3}\] but some times you can do it in your head

OpenStudy (anonymous):

is it y or f^-1(x)?

OpenStudy (anonymous):

\(f^{-1}\)

OpenStudy (anonymous):

okay, thanks!!